Reports of 57163
Environmental Monitoring (Environmental Monitoring)
Tringali, Fiori, Paoletti - 23:51 Monday 11 December 2023 (62679) Print this report
Hrec structures around 18 Hz

The Bruco page of Dec 7th shows coherence around 18 Hz between Hrec signal and environmental sensors: ENV_DT_ACC_Z and ENV_DT_MAG_V, installed on the top East flange and on the East tower base, respectively. As observed during the O3 run, the coherence at 18.55 Hz is due to the Supply air fan frequency of the CEB clean room air conditioning system.

About the coherence at ~18.77 Hz, we investigated the source of this line more than one year ago (#55963). The investigations highlighted that the line seems to be associated with a resonance mode of the DT tower or something inside it (safety structure). Indeed, Irene noticed that the amplitude of that line had been reduced during the DT tower venting in Nov 2020 (#49949).

Figure 1 shows a frequency zoom in the band (17-19) Hz of  the spectra and coherence between Hrec and the following sensors:

  • coh~0.8-0.9 @18.75Hz :  *DT_ACC*, *EDB_ACC*, *DT_MAG*, Sa_OB_F0_x  and NN_*_233 and 234 (geophones in DET-lab and in DET EE room, respectively) ;
  • coh~0.4-0.6 @18.55Hz  :  ENV_DT_ACC_Z, ENV_EDB_ACC_Z, Sa_OB_F0_x, ENV_DT_MAG* and NN_*_223 and 224;
  • coh~0.5-0.8 @17.5Hz, 18.02 Hz, 18.09Hz, 18.37Hz: significant coherence with the magnetometer on the tower base
  • coh~0.46 @(18-18.37) Hz: accelerometers  *DT_ACC* and *EDB*ACC* 

In Figure 2, it is observed coherence between the magnetometer *DT_MAG_V and the accelerometers on the flange, on EDB and geophones #233-234 at (18-18.77) Hz.

 Further investigations will follow.

 

Images attached to this report
AdV-INJ (Electro optical modulation system (EOM))
nocera - 23:24 Monday 11 December 2023 (62704) Print this report
RAMS Servo --preliminary tests

A. Masserot, F. Nocera, P. Spinicelli

Today we set off on the long journey that will bring one day to the commissioning and use of the RAMS Servo.
Today's goal was simple enough: verify whether or not it was possible to lock the ITF adding on the signal path the first stage of the (current version of) the RAMS Servo.
Details about the hardware and its characterization are in VIR-1061B-23.
After some initial legwork to have a reference to compare results with and to acquire familiarization with the setup to use (DAC reconfiguration etc., thanks to Alain), some adjustments, and a few mistakes, we finally found a configuration that allowed us to complete the first step, which was the measurement of the phase delay in the new configuration (Mixer Unit in, no BP filter).
Once that was done, we could compensate for it and try to lock the ITF, shortly berfore 20 local time.
The lock acquisition went smoohly and around 21.20 we reached Low Noise 3 with a good Sensitivity (around 36 Mpc).


At this point we wanted to attempt a first Line injection, but reloading the LNFS_Demod we caused un unlock do to the change in phase of the LNFS (the phaase delayed measured was not saved there)
After this, we started to relock.

Again, then lock acquisition went smooth for the second time.
Once in LN3, we started with the line injections. We chose two frequencies where the sensitivity seems quite clean: 121.1Hz and 721.0Hz.
Injection:
21.18.10 UTC @121.1Hz Amp 2e-3V
21.24.05 UTC @121.1Hz Amp 4e-3V
21.27.50 UTC @121.1Hz Amp 1e-2V

The injected line is clearly visible in Hrec.

then, we moved to the 721:

21.32.50 UTC @721.0Hz Amp 1e-3V and 121.1Hz OFF
21.35.40 UTC @721.0Hz Amp 0 ==> clean data
21.37.40 UTC @721.0Hz Amp 5e-3V
21.39.50 UTC @721.0Hz Amp 1e-2V
21.43.30 UTC @721.0Hz OFF

We then wanted to redo the injections with the SR_TY misaligned (i.e. best sensitivity at low frequency) but we unlocked when moving the SR for a human error.

In conclusion, the very first test can be considered successful. Weneed to repeat it misaligning the SR and possibly injecting white noise.
We leave the Mixer unit on the modulation signal path

To be noted: in order to keep a comparable modulation index for the 6MHz, we had to increase the setting on the LNFS up to 13dBm. This setting has been stored in the INJ_MAIN.INI file.

Detector Operation (Operations Report)
berni - 23:03 Monday 11 December 2023 (62702) Print this report
Operator Report - Afternoon shift

The planned activity on INJ_RAMS servo test was carried  on without issues.

The lock acquisition was succesfully tested a few times; the unlocks were due to the activity.

Activity concluded, ITF left in relocking to LOW_NOISE_3 with the autorelock engaged.

Images attached to this report
AdV-ISC (Commissioning up to first full interferometer lock)
bersanetti, mantovani - 18:48 Monday 11 December 2023 (62700) Print this report
Study of non-linear effects, new algos for OS, DCP and SR with DARM line at lower frequency

Today's shift was devoted to two activities:

  1. study of non-linear effects, by increasing particular injected lines and observing effects on hrec;
  2. update of the main ISC processes (LSC_Acl, LSC_Acl_moni, ASC_Acl and ASC_pre) with the introduction of a new DARM line at low frequency (33.7 Hz) to be used for a tentatively better signal for the OS/SRCL_SET loop; additionally, we wanted to get rid of the DARM_DCP_LF line at 176.3 Hz and replace it with the 74.4 Hz which is always on anyway; additionally, we wanted a new set of SR AA signals which use the same technique of the current one (DARM double demodulation and phi extraction), but demodulated at 74.4 Hz instead of 491.3 Hz.

Study of non-linear effects:

We started with the ITF in LOW_NOISE_3; we opened the SR loop and misaligned SR (+0.7 urad TY) to null the CMRF LF and maximize the BNS Range; then we injected back the SRCL line (off at that point), increasing its amplitude over several steps:

8:49 UTC 120 sec SRCL line ampl 0

8:52:0UTC 120 sec SRCL line ampl 0.001

8:56 UTC 120 sec SRCL line ampl 0.002

9:00 UTC 120 sec SRCL line ampl 0.004

9:03 UTC 120 sec SRCL line ampl 0.008

9:07 UTC 120 sec SRCL line ampl 0.016

9:10 UTC 120 sec SRCL line ampl 0.032

9:13 UTC 120 sec SRCL line ampl 0.05

Since the line is visible in hrec we could see already it induces other peaks (see Figure); more analysis will follow.

Then we closed again SR to restore its initial aligned position, then we misaligned again with the same purpose (although this time the best Range was achieved at +1.1 urad TY), then we started with the DIFFp TX line; given its low frequency (4.8 Hz) it is not visible directly in hrec and it will need offline analysis:

10:07:00 UTC SR AA open;

10:10 UTC to 10:32 UTC: SR misalignment (+1.1 urad TY) and ITF stabilization;

10:32 to 10:37 UTC: progressively increased DIFFp TX line amplitude;

10:37:00 UTC : 180 s with DIFFp TX line at 1e-3 amplitude, which is 10x the amplitude we have in LOW_NOISE_2 (in LOW_NOISE_3 is turned off completely).

Unfortunately we unlocked before moving to TY, so we moved to the second topic.


New DARM line at 33.7 Hz and restart of ISC processes:

We then moved to complete the work already started a while back, adding/changing a few things with respect to the code already deployed offline.

An initial disclaimer about nomenclature, which I think has become a bit obscure over time to some:

  • with "DARM line" we intend the main one, used for a number of things, which is injected at 74.4 Hz;
  • back in the day we added two lines to be used for the DCP computation, although this changed over time; these are called "DARM LF line" (176.3 Hz) and "DARM HF line" (491.3 Hz) as they were intended as lower than the DCP and higher than the DCP respectively; the former is now unused, not deleted, and kept to zero; the latter is used for many important things, and has been left unchanged by the work described here and in the referenced entry;
  • we introduce now a new "DARM LLF line", injected at 33.7 Hz; its purpose is to hopefully deliver a better signal for the OS, as it is closer to where the OS actually is.

In order not to have too many lines, as we introduced a new one we wanted to get rid of the DARM LF line, as it is used only for the DCP monitor, evaluated as the ratio of the mag of the demodulation of DARM at the HF and the LF line frequencies; now, instead, we replace the 176.3 Hz line with the 74.4 Hz one, which is always on as it is used also for the DARM UGF and other things.

So, this is the recap of the work done:

  • Optical Spring / SRCL_SET: with the new LLF line, we duplicated the whole OS code (_LLF_ version of the same channels), so now we have the same monitor as before, plus the new one, which is online but not calibrated just yet; for this reason, we changed a bit the code for the SRCL_SET servo, and now SRCL_SET_INPUT has become a weighted sum (ACL_SUM_CH object) of the two signals provided by the two machineries; right now, the usual signal has weight 1 and the new one weight 0; both of them have their calibration upstream with respect to such weight, and those would be the ones to be measured and set; a new set of demodulation filters at the LLF frequency have been added to LSC_Acl as well;
  • Double-Cavity Pole monitor: as explained above, now the LSC_DCP_moni_mad_cal channels is computed as the ratio of the magnitudes of the DARM demodulations at 491.Hz and 74.4 Hz, where the latter replaced the demodulation at 176.3 Hz; the calibration in CARM_NULL_1F (which was reasonably correct in Hz) has not changed considerably, at least at order zero; instead, the calibration in LOW_NOISE_2 (which was already incorrect in Hz, as it peaked at ~ 3000) has strongly changed, and now a high value of the DCP (cross-checked with the CAL estimator) corresponds to ~ 98 in those arbitrary units;
  • SR alignment signals: we added a new set of signals for the SR alignment; these use the same technique as the current ones, which means a double demodulation of DARM at a DARM line (phi extraction) and then at a SR TX/TY line; the DARM demodulation is done at 74.4 Hz (instead of 491.3 Hz) and as such they should be more sensitive to the DARM optical gain rather then the DCP; these signals are named SR_T*_DCP_phi_B1_I and are generated in ASC_pre as the other ones; for internal consistency, though, once they are acquired in ASC_Acl they get the name the current one have, and the current ones have an additional _HF_ segment in their name to reflect the HF demodulation instead of the 74.4 Hz one. This should be transparent, though, as those duplicated were and are not saved in the DAQ, and the aforementioned channels from ASC_pre should be used instead, as usual. The new channels are online but they are still missing demodulation phase and calibration.

After the restart of the processes, we verified that the OS demodulation phase was still good, we retuned the demodulation phases for the SR alignment signals (both the B1p- and the B1-based ones), and we relocked up to LOW_NOISE_2.

The Automation has been already update to reflect all the changes reported above; in particular:

  • the new DARM LLF line is reset in DOWN, and it is turned on during lock acquisition; its amplitude is almost double the one of the 74.4 Hz one in CARM_NULL_1F, and equal to that in later states; differently from the 74.4 Hz line, it is not turned on before CARM_NULL_1F;
  • duplicated commands for the new OS monitor have been changed, numerical values still need to be updated, following calibration;
  • the sensing matrix for the SR alignment already reflects the changes in the SR signals, and a new section will need to be uncommented (commenting the current one) in case of use; numerical parameters are there but will need measurements.
Images attached to this report
Detector Characterisation (Glitches)
Paoletti, Fiori, Tringali, Direnzo - 17:27 Monday 11 December 2023 (62701) Print this report
Comment to Hrec glitches after the power outage (62576)

Today there was a long lock that lasted more than 12 hours.The ITF unlocked after 10:41 UTC.

The long and annoying series of glitches in Hrec (well detectable in the DQ_BRMSMonHrec_BRMS_HREC_HOFT_FREQ_BAND_85_95_Hrec_hoft_16384Hz channel) still had a time interval of 1320s +/- 30s (i.e. 22m), apart from the last two before unlocking which had a strange, delayed value of about 26m.

The ITF relocked at 14:36 UTC and again we had another couple of glitches with an interval of 1620s (27m), while the third one was back to the initial 1320s (22m). Then again a delay of 1511s (25m).

It is to be figured out what could be the reason for this variations in the time interval, which had been only shortening in recent months instead. No particular action seems to have been performed on the ITF or Virgo's environment. Let's see what happens from now on.

Images attached to this comment
Detector Operation (Operations Report)
gherardini - 15:54 Monday 11 December 2023 (62697) Print this report
Operator Report - Morning shift
This morning I found the ITF locked in LOW_NOISE_3 state, it unlocked after more then 14 hours of lock; the foreseen ISC activity, carried out by Maddalena and Diego, went on all the shift without any major problem; activity completed at the end of the shift and ITF relocked at LOW_NOISE_2.
Images attached to this report
AdV-PSL (Laser Bench setup)
chaibi, coulon, gosselin - 11:06 Monday 11 December 2023 (62699) Print this report
Investigations on the IMC loop
During the week end, we investigated the IMC loop behavior in order to improve it:
- Measurement of the IMC loop: UGF 80 kHz, phase margin 19° (22dB of attenuation); UGF 100 kHz, phase margin 17° (19dB of attenuation)
- Effect of the SL electronic loop: the gain of SL electronic loop was increased then decreased by 4dB, no effect observed on the possibility to improve the IMC loop
- Measurement of the pEOM transfer function: the IMC was locked in 1/f regime and a perturbation signal was injected into (i) the pEOM preamplifer (ii) directly the pEOM. Their behavior was normal.
- Measurement of the IMC_ref photodiode response : The IMC was unlocked and the its end mirror tilted. The pahse-pEOM was swapped by an amplitude-pEOM. The response of the pEOM was measured by using calibrated photodiode on the laser bench ; then the response of several photodiode on the LB and EIB (to measure the TF of SL and PMC) including the IMC_ref were measured -> Normal behavior.
- Measurement of the SL optical injection frequency band : 1.9MHz. After increasing the power injected into the SL (20% recovered from the fiber amplifier pick-off beam) the injection frequency band was increased by 10% as expected. This allowed to improve the IMC loop : UGF 115 kHz, phase margin 19° (18dB attenuation). This effect needs to be investigated.

Once the data fully analyzed, plots and more accurate numbers will be provided in a separate logbook entry.
Detector Operation (Operations Report)
gherardini - 8:01 Monday 11 December 2023 (62696) Print this report
Operator Report - Initial status
ITF State: LOW_NOISE_3
Quick Summary: all suspension loops closed; all SBEs loops closed;
No activity ongoing.
Images attached to this report
AdV-SGD (FDS commissioning)
vardaro - 0:16 Monday 11 December 2023 (62695) Print this report
Comment to FIS Injection with SR Misaligned (62684)

The retroreflector is removed.

I did manually FMODERR (sidebands are off)

Then I unclipped the IR in the faraday.

I left the system in FLT_IR_AA_FMODERR_TUNED

Images attached to this comment
Detector Operation (Operations Report)
menzione - 22:59 Sunday 10 December 2023 (62693) Print this report
Operator Report - Afternoon shift

Upon arrival I found ITF LOCKED_IR and INJ/PSL activity on noise hunting (Chaibi, Coulon, Gosselin) still in progress.
It has been concluded at 15:30 UTC. Then I started to recover ITF. I realigned PR and SR in ACQUIRE_DRMI.
During the lock, at 15:38 UTC,  I had to close (open and close) B2 QD2 galvo loop.
Upon fully recovered ITF up to a stable LN3 I started the calibrations not performed yesterday eveving.

Planned Calibrations
- 16:15 UTC - Measure again checkHrec and optical response and sensitivity [failed due to unlock after 28'' at 16:44 UTC]
- 16:45 UTC - Measurement of the TF between CAL and Sc channels [ok]
- 16:50 UTC - Measurement of actuators response for PR and BS mirrors [ok at 2nd attempt]
- 17:45 UTC - Measurement of actuators response for SR mirror [failed twice]
- 17:48 UTC - Measurement of actuators response for NI,WI mirrors [ok]
- 19:12 UTC - CALIBRATED_DF_DAILY [failed due to unlock after 12'' at 19:24 UTC]
- 19:26 UTC - 5 minutes of NI Single Bounce [ok]
- 19:32 UTC - 5 minutes of WI Single Bounce [ok]

ITF left in LN3 (AUTORELOCK_FAILSAFE)

Sub-system reports

ISYS
Under request of Gosselin I lowered the threshold of INF_EER_CHILLER_FLOW from 5 to 4.9.

Images attached to this report
Detector Operation (Operations Report)
menzione - 22:43 Sunday 10 December 2023 (62694) Print this report
Comment to Operator Report - Initial Status (62685)

concerning the TCS NI Chiller flow decreasing, Yesterday at 14:00 UTC I went in TCS chiller room to check the status of the NI aux pipe, I noticed that it was a bit strangled. I stretched it out in a correct position and the flow came back to the normal value. (plot attached)

Images attached to this comment
Detector Operation (Operations Report)
amagazzu - 16:01 Sunday 10 December 2023 (62692) Print this report
Operator Report - Morning shift

The shift was dedicated to the planned INJ/PSL, noise hunting, carried out by Chaibi, Coulon, Gosselin.
Upon reaching the site, I found the ITF in LOW_NOISE_2 since yesterday night, I was able to reach LOW_NOISE_3 at 7:07 UTC. At 7:24 UTC I launched the daily Calibration, but the ITF unlocked at 7:37 UTC.
At 8:19 UTC, under request of the crew, I set the ITF in DOWN state to allow the planned work. At 12:35 UTC the activity in Laser Lab was concluded, but not all loops have been restored. 
Activity still in progress, ITF left in LOCKED_ARMS_IR.

Other activities carried out during the shift:
- Fd package update (see report #62691);

Calibrations
7:24 UTC - CALIBRATED_DF_DAILY, ITF unlocked at 7:37 UTC;
7:44 UTC - 5 minutes of NI Single Bounce;
7:50 UTC - 5 minutes of WI Single Bounce;

Images attached to this report
AdV-DAQ (Data collection)
masserot, mours - 11:40 Sunday 10 December 2023 (62691) Print this report
Comment to SDB2_Fb stopped(crashed) (62676)

The issue has been fixed in the Fd package v8r60p3 . A new release of the TolmFrameBuilder package v17r14p7 has been built and deployed on all the TolmFrameBuilder's server .

Operations performed between 2023-12-10-09h30m27-UTC and 2023-12-10-09h41m24-UTC

The attached plot (time as LT not UTC) show the memory usage on the SDB2_Fb's rtpc (rtpc5) according the Fd version: this TolmFrameBuilder was running the lastest release since 2023-12-09-11h54m13-UTC

Images attached to this comment
Detector Operation (Operations Report)
amagazzu - 8:11 Sunday 10 December 2023 (62690) Print this report
Operator Report - Initial Status

ITF State: LOW_NOISE_2
ITF Mode: LOCKED (AUTORELOCK_FAILSAFE Engaged)
Quick Summary: ISYS in standard state; All SBE Loops closed; All Suspensions Loops closed; INJ LaserChiller Laser Flow decreasing;
Activities ongoing since this morning: Daily Calibrations;

Images attached to this report
Detector Operation (Operations Report)
menzione - 23:00 Saturday 09 December 2023 (62689) Print this report
Operator Report - Afternoon shift

Upon arrival I found ITF DOWN and INJ/PSL activity on noise hunting (Chaibi, Coulon, Gosselin) still in progress.
Meanwhile Bulten recovered, from remote, the actuator horizontal 0 for the EIB-SAS.
It has been concluded at 17:30 UTC. Then I started to recover ITF. I realigned both cavities (very far), PR and SR in ACQUIRE_DRMI.
Upon fully recovered ITF up to a stable LN2 I started the planned calibrations.

Planned Calibrations:
-  9:19 UTC - Measurement of the TF between CAL and Sc channels [ok]
- 19:25 UTC - Measurement of actuators response for PR and BS mirrors [ok]
- 20:15 UTC - Measurement of actuators response for SR mirror [failed twice]
- 20:22 UTC - Measurement of actuators response for NI,WI mirrors [ok]
- 21:30 UTC - Measurement of actuators response for NE,WE actuators and NI,WI,BS marionnettes  [ok]

At 22:00 UTC Calibrations interrupted, ITF left in LN2 (AUTORELOCK_FAILSAFE) - No time to perform other calibrations.

Images attached to this report
AdV-SBE (EIB-SAS commissioning)
bulten - 16:28 Saturday 09 December 2023 (62688) Print this report
EIB SAS balancing
The actuator horizontal 0 for the EIB-SAS was above 7 V and gave orange alarm. I tried to remedy that now, since there was a slot in the commissioning with the beam unlocked.
As I suspected, (see report 62368) the stepper motors pulled the bench to the endswitch; that is because they have not been used since the power failure in November. Unfortunately, the
horizontal microsprings are al very close towards the end; the Ty rotation of the bench is out of reach of the microsprings. I reduced the actuation of horizontal actuator 0 from
7 to 5.5 Volt, but that could only be reached by having higher voltages (3 on actuator 2 and 4.5 on actuator 1).

The only way to get the zero set point in range of the bench is by re-doing the cabling that pulls on the external injection bench and introduces torsion. This operation has been tried
already twice (last report logbook 60754); at that time we could reduce the Ty offset from -1500 to -600 microrad and the actuator voltage close to zero.
However, that was the furthers that the cables on top of the bench could be pulled.

I propose to continue running in the current condition; the actuator voltages are somewhat high but the noise on the bench is so high that the actuator noise plays zero role.
I think that we could rebalance the bench by shifting countermass weights of maybe by mechanically changing the tilt stabilization a bit; this is however a delicate operation and
would require my presence (or Alessandro); I prefer to do such an intervention at a moment that we can simultaneously fix the SQB2 and SIB2 stuck fishing-rods on filter0.
Detector Operation (Operations Report)
amagazzu - 16:01 Saturday 09 December 2023 (62687) Print this report
Operator Report - Morning shift

The shift was dedicated to the planned INJ/PSL, noise hunting, carried out by Chaibi, Coulon, Gosselin.
Upon reaching the site, I found the ITF in LOW_NOISE_3 since yesterday night. At 7:04 UTC I launched the daily Calibration, but the ITF unlocked at 7:24 UTC.
At 7:55 UTC, under request of the crew, I set the ITF in DOWN state to allow the planned work. Activity still in progress.

Calibrations
7:04 UTC - CALIBRATED_DF_DAILY, ITF unlocked at 7:24 UTC;
7:30 UTC - 5 minutes of NI Single Bounce;
7:36 UTC - 5 minutes of WI Single Bounce;
7:54 UTC - Lock of CITF, unlocked after a minute.

Sub-system reports

SQZ
From 10:48 UTC to 11:29 UTC, under request of Sorrentino, I closed the Squeezing SQB1 Shutter.

Images attached to this report
AdV-SGD (FDS commissioning)
vardaro - 14:35 Saturday 09 December 2023 (62686) Print this report
Comment to FIS Injection with SR Misaligned (62684)

I rotated the SQB1_HWP2 to send back SQZ to HD detector

I measured 5.85 dB of SQZ and 8.35 dB of ASQZ. With 8.8 dB of generated squeezing.

I asked to Henning to switch back in BAB. The SQB1 retroreflector is still on.

Detector Operation (Operations Report)
amagazzu - 8:24 Saturday 09 December 2023 (62685) Print this report
Operator Report - Initial Status

ITF State: LOW_NOISE_3
ITF Mode: LOCKED (AUTORELOCK_FAILSAFE Engaged)
Quick Summary: ISYS in standard state; All Suspensions Loops closed; EIB act0 close to saturation, all other SBE loops close and in stanrda condition; INJ LaserChiller Laser Flow decreasing; TCS NI Chiller flow decreasing;
Activities ongoing since this morning: Daily Calibrations;

Images attached to this report
Comments to this report:
menzione - 22:43 Sunday 10 December 2023 (62694) Print this report

concerning the TCS NI Chiller flow decreasing, Yesterday at 14:00 UTC I went in TCS chiller room to check the status of the NI aux pipe, I noticed that it was a bit strangled. I stretched it out in a correct position and the flow came back to the normal value. (plot attached)

Images attached to this comment
AdV-SGD (FDS commissioning)
vardaro, bonavena, delaurentis, zendri - 23:03 Friday 08 December 2023 (62684) Print this report
FIS Injection with SR Misaligned

This is a brief entry the full one will follow in the next days.

Today we injected SQZ in normal configuration: SR aligned ando also with SR Misaligned by +1.1urad in TY.

The SR misalignment was completed at 17:22 UTC. Then we started the characterization with Injections and a phase scan.

At 18:43 we started a long run of SQZ injected with SR misaligned

At 18:50 we put the ITF in LN3

At 19:00 UTC we put the ITF STATUS node in autorelock filesafe

The maximum range reached during the injection was 48.8 Mpc.

 

We left the system in the following way:

  • ITF_LOCK: LOW_NOISE_3
  • ITF_STATUS: Autorelock filesafe
  • SQZ_MAIN: SQZ_INJECTED
  • SQB1_Shutter: Open
  • SR_TY: Misaligned by +1.1urad 
Images attached to this report
Comments to this report:
vardaro - 14:35 Saturday 09 December 2023 (62686) Print this report

I rotated the SQB1_HWP2 to send back SQZ to HD detector

I measured 5.85 dB of SQZ and 8.35 dB of ASQZ. With 8.8 dB of generated squeezing.

I asked to Henning to switch back in BAB. The SQB1 retroreflector is still on.

vardaro - 0:16 Monday 11 December 2023 (62695) Print this report

The retroreflector is removed.

I did manually FMODERR (sidebands are off)

Then I unclipped the IR in the faraday.

I left the system in FLT_IR_AA_FMODERR_TUNED

Images attached to this comment
AdV-NNC (CEB)
Tringali - 20:39 Friday 08 December 2023 (62683) Print this report
CEB-LEV2 restarted

Today at 12:47:45 UTC, I restarted the acquisition process of the CEB-LEV2 array which had been interrupted since November 30th at 15:27 UTC.

Detector Operation (Operations Report)
rolland - 18:13 Friday 08 December 2023 (62682) Print this report
Comment to Operator Report - Morning shift (62631)

Some results of "daily calibration" from December 1st, 4th and 8th

Injections to check hrec (estimation of bias)

The first four plots show, for a given actuator to inject hinj, the ratio hrec/hinj for different datasets, which are rather stable over the 3 different days.

Then, these dataset have been averaged (even if not everywhere compatible with statistics, mainly for f<30 Hz): see three examples in figures 5, 6 (for MirNE) and 7 (for PcalNE).

The 8th figure shows the superposition of the averaged results for the different actuators: they all see the same bias of h(t), with currently ~2% maximum differences in absolute amplitude.

 

Measurement of optical responses

The model for DARM sensing has been updated, and used for the analysis of some optical response measurements.

The model fitted on these data is based on the model proposed by Michal in logbook entry 62250.

A preliminary results are shown in figure 9 for optical response to a motion of NE mirror. Ann optical spring has been fixed with frequency of 4 Hz and Q = 10 (which is not exact looking the low frequency). The fitted parameters, for data from December 1st, are:

  • optical gain 1.6e9 W/m
  • frequency of the zero: 368.7 Hz
  • SR detuning angle: ~0.                    which gives a DCP at 368.7 Hz with Q = 0.50
  • extra-delay of -13 µs       -> this being quite close to 0 gives confidence in the data modeling and fit result.

The optical response when moving the WE mirror few minutes apparts is shown in figure 10, with the following parameters:

  • optical gain 1.7e9 W/m
  • frequency of the zero: 375.0 Hz
  • SR detuning angle: ~-7 deg.                    which gives a DCP at 374.3 Hz with Q = 0.502
  • extra-delay of -18 µs

More data from 4th and 8th December are still to be analysed to estimate if this fit is stable and can highlight optical response different from a simple pole (i.e. SR detuning).

 

NOTE: in the last two "daily calibration" (4 and 8 December), the ITF unlocked during the injections for the optical measurements (injections on DARM and on the different mirror actuators), but not always at the same injection.

Images attached to this comment
Detector Operation (Operations Report)
Sposito - 16:35 Friday 08 December 2023 (62678) Print this report
Operator Report - Daily shift

Today's shift was dedicated to SGD - squeezing injection with misaligned SR. Activities went on as planned, and we experienced some unlocks. The shift went on without any particular difficulties. Activities are still in progress. 

Sub-system reports

DET
07:42 UTC CALIBRATED_DF_DAILY failed

Images attached to this report
AdV-DAQ (Data collection)
masserot, mours - 15:58 Friday 08 December 2023 (62681) Print this report
Comment to SDB2_Fb stopped(crashed) (62676)

The crash of this night’s TolmFrameBuilder SDB2_Fb is due to a lack of memory due to a memory leak (plot)  present since April 2023 (plot).

After some investigations, the memory leak seems present on all the TolmFrameBuilders building a frame at 5Hz and filtering at 1Hz  the pressure and temperature channels of the in air box  photodiodes.

A new Fd's version will be delivered to fix this issue

Images attached to this comment
AdV-COM (1/√f noise)
mwas - 14:49 Friday 08 December 2023 (62680) Print this report
Optical gain and 1/sqrt(f) noise over past month

Figure 1 shows the optical gain at ~60Hz since Oct 25. In blue locks in LN2 and in red in LN3. There are a few spikes in gain during the SR misalignment tests. Overal for the aligned SR case there has been a nice increase in optical gain of ~20% since end of October.

Figure 2 There has been also a step in the SR+arm cavity pole frequency from 350Hz to 400Hz between Dec 4 and Dec 5, and it doesn't seem directly related to a change in optical gain at low frequency.

Figure 3 on this timescale of 1 month, the BNS range upper limit seems vaguely proportional to the optical gain.

Figure 4 shows that during that time the amount of noise on B1 between 100Hz and 150Hz has not changed significantly

Figure 5 shows that the h(t) noise floor has changed much more than the B1 noise. Which confirms that the improvement in range, is due to the increase in optical gain with a fixed background noise in terms of mW/rtHz on B1.

Figure 6 note that the height of the calibration lines in h(t) has changed, because the amplitude of the calibration line correction has been changed over this time period.

My conclusion from this is that the TCS tuning is working, and the goal of it should be to increase the optical gain at ~60Hz (V1:Hrec_ORgain_V1:Sc_NE_MIR_Z_CORR_62.5) when SR is aligned. And then we can use SR misalignment to increase the optical gain further.

Images attached to this report
Search Help
×

Warning

×